OpenStudy (anonymous):

NEED HELP PLEASEEE!!! I WILL GIVE MEDALS!!! What is the value of the variable A that makes the statement true? A + (–7) + 3 = 0 A. –4 B. –3 C. 4 D. 7

OpenStudy (nikato):

What is (-7)+3?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@nikato -4

OpenStudy (nikato):

Yes So A-4=0 Now add 4 to both sides so you're only left with A on the left side

OpenStudy (anonymous):

I'm very confused. Can I see an example?

OpenStudy (anonymous):

A-4=0 +4=0+4 which equals A=4

OpenStudy (anonymous):

You 100% sure?

OpenStudy (nikato):

Yea. That's right!

OpenStudy (nikato):

A-4=0 +4. +4 --------- A-4+4 = 0+4. Remember -4+4=0 So A+0= 0+4 Simplify A=4

OpenStudy (anonymous):

yesss. medal?